There are two homes. Each has a rectangular garden,with a perimeter of 24 yards. One garden has an area that is 8 square yards more than the other . What are the areas of the two gardens?
Answer by richwmiller(17219)

2(L+W)=24
L+W=12
1,11=11
2,10=20
3,9=27
4,8=32
5,7=35
6,6=36
l*w=L*W+8
L+w=12
L+W=12
3*9=27 area one
5*7=35 area two
35-27=8